OpenStudy (noodya123):

PLEASE HELP A certain forest covers an area of 3000km^2. Suppose that each year this area decreases by 5.75% . What will the area be after 7 years? round your answer to the nearest square kilometer.

OpenStudy (flamingfrost25):

5.75*7=??? then you take that and take the decimal place over 2 spaces to the left and subtract from your 3000

OpenStudy (flamingfrost25):

how so

OpenStudy (mathstudent55):

Actually, I'm not sure. This may or may not be exponential decay.

OpenStudy (mathstudent55):

If this is exponential decay, that means the first year, you loose 5.75% of the original area of 3000 km^2. When you start the second year, you have 5.75% of 3000 km^2 less than the original area. Then when you lose 5.75% of the new smaller area in the second year, you are not loosing 5.75% of the original 3000km^2. You are loosing 5.75% of the remaining area after having lost the original loss of the first year.

OpenStudy (mathstudent55):

The formula for exponential decay is: \(F = P(1 - r)^t\)

OpenStudy (noodya123):

i got the answer it was 1982 thank you everyone tho

OpenStudy (mathstudent55):

\(F = 3000(1 = 0.0575)^7\) \(F = 1981.94\)